OpenStudy (anonymous):

Write a recursive definition for the sequence 11,8,5,2

OpenStudy (ybarrap):

$$ a_0=2\\ a_1=5=a_0+3\\ a_2=8=a_1+3\\ a_3=11=a_2+3 $$ Do you see the pattern? So, the nth number can be found from the (n-1)st number using the formula $$ a_0=2\\ a_i=a_{i-1}+3 $$ But we want to go backwards from 11,8, to 2 not 2 to 11, so $$ a_0=11\\ a_1=8=a_0-3\\ a_2=5=a_1-3\\ a_3=2=a_2-3 $$ So, the nth number can be found from the (n-1)st number using the formula $$ a_0=11\\ a_i=a_{i-1}-3 $$ Which is the recursive equation we sought. Make sense?

OpenStudy (anonymous):

yes sir is does thank you:)